how to find the variance and standard deviation of a discrete random variable

1
Expert's answer
2022-02-28T15:33:42-0500

Let X be a discrete random variable, such that

"P(X=x_1)=p_1"

"P(X=x_2)=p_2"

...

"P(X=x_n)=p_n"

Then its variance can be found as "V(X)=E(X^2)-E^2(X)" , where E(X), E("X^2") - first and second central moment respectively, so

"E(X)=\displaystyle\sum_{i=1}^np_i*x_i"

"E(X^2 )=\displaystyle\sum_{i=1}^np_i*x^2_i"

Discrete random variable can take infinite(countable) amount of values, in that case sum will be from 1 to infinity, and the point is to find the sum of the infinite row

Standard deviation can be found as a square root of the variance, so

"\sigma(X)=\sqrt{V(X)}"